OpenStudy (anonymous):

1/(a-b) + 4/(b-a) - 8/(a+b) - (11a - 5b)/(b^2-a^2) simplify and plz explain your answer...

OpenStudy (anonymous):

note (a-b) = -(b-a) , and (b^2 - a^2) = (b+a)(b-a) Multiply through by whatever you need to to get it all over this fraction and see what cancels. I can;t be bothered with the leg work

OpenStudy (anonymous):

Oh, it cancel down very nicely. You could say it cancels down to virtually 'nothing'...

OpenStudy (anonymous):

so you mean it's zero?

OpenStudy (anonymous):

I'm afraid you'll have to work it out first

OpenStudy (anonymous):

I'll give you a clue, 1/(a-b) = -1/(b-a) so the first two become 3/(b-a) If you want it all over b^2-a^2 you need to multiply this by (b+a) and the other fraction over (a+b) by (b-a)

OpenStudy (anonymous):

I thought you said you were gonna test my skills. I am disappoint. Goodnight.

OpenStudy (anonymous):

well geez I was looking for help not criticism...

OpenStudy (anonymous):

goodnight yourself